A ball is thrown vertically upwards with a velocity of 20 m/s from the top of the building of height 25m from the ground (i).How high will the ball reach? (ii)How long will it take to reach the A ground?


1
Expert's answer
2020-12-25T14:05:41-0500

As per the given question,

Initial speed of the ball (u)=20 m/s

Height of the building (h)= 25m

At the top most point the speed of the ball become zero hence,

Let the height reached by the ball from the top of the building be H and gravitational acceleration be "g=10 m\/s^2" .

applying newton's third equation of motion,

"v^2=u^2-2gH"

Now, substituting the values,

"\Rightarrow 0=20^2-2\times g\times H"

"\Rightarrow H = \frac{400}{2\times 10}=20 m"

Hence, total height from the ground be "=H+h = 25+20=45m"

Let the time taken by the ball to reach the ground be t.

"-h=ut-\frac{gt^2}{2}"

"\Rightarrow -25=20t -5t^2"

"\Rightarrow 5t^2-20t-25=0"

"\Rightarrow t^2-4t-5=0"

"\Rightarrow t^2-5t+t-5=0"

"\Rightarrow t(t-5)+1(t-5)=0"

"(t-5)(t+1)=0"

Hence time taken by the ball to reach the ground be 5 second.